#22
Elasticity For each of the following demand functions, find (a) E, and (b) values of q (if any) at which total revenue is maximized. 20. 25,000 50p 22-9 = 48,000-1 24. q10 In p e 19. q 50- 4 0p2 be 21. 37,500 5p2 23. p 400e 0.24
Answer #1

450 - 1 (a) Elasticity # - ya 200-P (b) TR is max when E l E -200- 2) 200-P=P 2p-200 p 100 니 20 = 25000-50P (a) E= - ΧΡ x? =-2s,000 SDP 25,000-50 (250%) Aj = 25,000-12500e → 12,S00 = TR ls max. at A,-12,500dP 31,500-Sp b) TR mnax.at E 37,soo -sp 37,soo ISP2 231, S00 2500 IS SD 37,soo 12,Soo = 2S,000 TRE nox, at p.25.co0コ NDTE : C